OpenStudy (anonymous):

[10.01] Which is the simplified form of 9 times x minus 18 all over 4 times x minus 8.? 9 times x minus 1 all over 4 times x minus 1. 9 over 4. 9 over 2. negative 9 over 4.

OpenStudy (akashdeepdeb):

9 times x minus 18 over four times x minus 8 can be written as \[\frac{9x-18}{4x-8}\] You can take 9 common from the numerator and 4 commonn from the denominator and use the reverse of the distributive property of arithmetic. Like so.. \[\frac{9(x-2)}{4(x-2)}\] You can then cancel \((x-2)\) thereby leaving you with 9/4 or 2.25. Understood? :)

OpenStudy (anonymous):

yea and i got 2.25 lol i understood it now

OpenStudy (anonymous):

Be D lol

OpenStudy (anonymous):

Ok, lets look at it in clear manner. (9x-18)/(4x-8) When you look at this you can see that you can actually factor out in both denominator and numerator. For the numerator you can factor out 9 so, 9*x-9*2 = 9(x-2) In denominator, 4*x-4*2 = 4(x-2) Now let's rewrite the fraction 9(x-2) / 4(x-2) Yes, you can cancel out the (x-2); so the answer is 9/4 We can cancel out (x-2) because let's say x=4, you will get 2 for (x-2) in both cases and you can cancel them out. Hope it helped.